A luxury car manufacturer keeps detailed records on all current owners of their vehicles. The marketing director takes a random sample of 9 customer records and finds an average age of 53.7 and a standard deviation of 6.5 for owners.

Assuming all conditions of inference were met, what is the lower limit for a 95% confidence interval for the average age of all owners of this luxury car?
Answer #1

Solution :

t /2,df = 2.306

Margin of error = E = t/2,df * (s /n)

= 2.306 * (6.5 / 9)

Margin of error = E = 5.0

Lower limit = 53.7 - 5.0 = 48.7